the mass in grams of HCI needed to react with 12.8g of alumimum



1
Expert's answer
2022-04-15T14:44:03-0400

"6HCl+2Al \rightarrow 2AlCl_3+3H_2"

M(Al)=27 g/mol

n(Al)=12.8/27=0.474 mol

n(HCl)=3n(Al)=3(0.474)=1.42 mol

M(HCl)=1+35.5=36.5 g/mol

m(HCl)=36.5 x 1.42=51.91 g